9. -/8 points OSColPhys2016 2.7.P.041 My Notes Ask Your Teacher Assume air resistance is negligible unless otherwise stated Calculate the displacement and velocity (in m and m/s, respectively) at the following times for a ball thrown straight up with an initial velocity of 12.7 m/s. Take the point of release to be Yo0 (a) 0.500s 1 m/s (b) 1.00s m/s (c) 1.50s m/s (d) 2.00s m/s Additional Materials Reading
Answer #1

(a) y1 = y0 + vit + 12at2

y1 = 0 + 12.7*0.5 - 0.5*9.8*0.52

y1 = 5.125 m

v1 = vo + at

v1 = 12.7 - 9.8*0.5

v1 = 7.8 m/s

--------------------------------------------------------------

In same way solve for other times

Part (b)

y2 = 0 + 12.7*1 - 0.5*9.8*1

y2 = 7.8 m

v2 = 2.9 m/s

Part (c)

y3 = 0 + 12.7*1.5 - 0.5*9.8*1.52

y3 = 8.025 m

v3 = 12.7 - 9.8*1.5

v3 = -2 m/s

Part (d)

y4 = 0 + 12.7*2 - 0.5*9.8*4

y4 = 5.8 m

v4 = 12.7 - 9.8*2

v4 = -6.9 m/s